OpenStudy (anonymous):

Find two real numbers that you are reasonably sure are not algebraic. Give your reason for suspecting that the value is not algebraic.

OpenStudy (anonymous):

I would say prime numbers simply because there isn't an algebraic equation to express them? That would just be my guess though...

OpenStudy (anonymous):

Ah, good thought. That does make sense I will start with that.

OpenStudy (zarkon):

prime numbers (in fact all integers) are algebraic.

OpenStudy (anonymous):

Is my prof asking a trick question then?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

Would numbers like pi and e be real but not algebraic?

OpenStudy (anonymous):

1 isn't algebraic?

OpenStudy (zarkon):

\(\pi\) and \(e\) are not algebraic

OpenStudy (anonymous):

really? prove it!

OpenStudy (zarkon):

1 is algebraic

OpenStudy (zarkon):

all integers are algebraic

OpenStudy (zarkon):

the proofs are non trivial...pi being harder than e

OpenStudy (anonymous):

yeah i was kidding

OpenStudy (zarkon):

I know ;)

OpenStudy (anonymous):

matter of fact i don't really remember the one for \(\pi\)

OpenStudy (zarkon):

I'd have to look it up...I saw it once at a conference

OpenStudy (zarkon):

i believe that \[\sum_{k=1}^\infty 10^{-k!}\] is transcendental too

OpenStudy (zarkon):

yep... http://en.wikipedia.org/wiki/Transcendental_number#History

OpenStudy (anonymous):

what about \(.1234567891011121314...\)

OpenStudy (anonymous):

@daniellerae9 is it clear what "transcendental" means in this context?

OpenStudy (anonymous):

Yes. I just learned what it meant the other day and forgot, transcendental numbers are just real numbers that aren't algebraic.